OpenStudy (anonymous):

Three vector a, b, c are given. Proove that if a┴b, a┴c, b┴c (┴ is symbol for perpendicular) and |a|=|b|=|c|=1 that a=b x a b=c x a c=a x b (it's vector product)

OpenStudy (anonymous):

Hey, I think some problem is there in the first value, the value of a.... It should be symmetrically a= b cross c.... Only then u can get the desired... And all vectors are following symmetry

OpenStudy (anonymous):

@saloniiigupta95 yeah, sorry. It should be: a=b x c b=c x a c=a x b

OpenStudy (amistre64):

hmm, would the formula:\[sin~\theta=\frac{u\times v}{|u|~|v|}\]be applicable?

OpenStudy (anonymous):

but that's definition of vector product....

OpenStudy (anonymous):

@amistre64 do you mean \[\sin \theta = \frac{ |a \times b |}{ |a||b| }\] ? With this formula I could show that the length of cross product is 1. Would that be enough to of a proof that a=b x c ? Perhaps I need to show that (b x c) and a are facing the same side too?

OpenStudy (amistre64):

yeah, thats the one i was thinking of :)

OpenStudy (amistre64):

sin(t) = sqrt(1-cos^2(t)) cos(t) = u.v/|u||v| sin(t) = sqrt(1-u.v/|u||v|); but it is given that u and v are parallel; so u.v = 0 sin(t) = 1 was something that i was thinking about